# Preserving Expert-Level Privacy in Offline Reinforcement Learning

### v1

@@ 1 INTRODUCTION
Figure 1: Training pipeline for Expert-level Differentially Private Offline RL given an expert set $\Pi=\{\pi_{1},\pi_{2},\dots,\pi_{m}\}$ and input trajectories logged into an offline dataset $D_{\Pi}$. $\mathcal{A}_{DR}$ (Algorithm 2) splits input trajectories and adds stable prefixes to $D^{\Pi}_{stable}$ discarding the rest to $D^{\Pi}_{unst}$. These are used to train any off-the-shelf offline RL algorithm as described in Algorithm 3 and learn an expert-level private policy $\pi_{private}$.
Leveraging existing offline datasets to learn high-quality decision policies via offline Reinforcement Learning (RL) is a critical requirement in many settings ranging from robotics (Fu et al.,, 2020; Levine et al.,, 2020) and recommendation systems (Bottou et al.,, 2013; Ie et al.,, 2019; Cai et al.,, 2017) to healthcare applications (Oberst and Sontag,, 2019; Tang et al.,, 2022). Correspondingly, there is now a rich literature on effective techniques (Fujimoto et al.,, 2019; Kumar et al.,, 2019, 2020; Cheng et al.,, 2022) for learning from such offline datasets, collected using one or more behaviour policies. An often overlooked aspect of this literature is, however, that in privacy sensitive domains such as personalized retrieval, advertising, and healthcare, the behaviour policies might reveal private information about the preferences / strategies used by the corresponding experts (users, advertisers, health care providers) whose decisions underlie the offline data. In such scenarios, we ideally seek to uncover the broadly beneficial strategies employed in making decisions by a large cohort of experts whose demonstrations the offline data is collected with, while not leaking the private information of any given expert. We call this novel setting offline RL with expert-level privacy, and design algorithms with strong privacy guarantees for it.
There has been a growing interest in differentially private RL, both in online (Wang and Hegde,, 2019; Qiao and Wang,, 2023; Zhou,, 2022; Liao et al.,, 2023) and offline (Qiao and Wang,, 2024) settings. The offline RL literature, which is most relevant to this work, primarily focuses on protecting privacy at the trajectory-level. However, protecting the privacy of an expert, who might contribute multiple trajectories to the demonstration dataset, requires significantly more careful techniques. Secondly, the prior works mostly focus on tabular or linear settings (Qiao and Wang,, 2023, 2024; Zhou,, 2022), where the offline RL problem can be effectively solved using count or linear regression based techniques, and is hence amenable to existing techniques for privatizing counts or linear regression models. In contrast, we make no assumptions on the size of the state space, the parameterization of the learned policy or the policy underlying expert demonstrations.
@@
  • We identify and formalize the question of offline RL with expert-level privacy and motivate it with practical examples. We also explain the inadequacy of prior algorithms in this setting.
  • We provide practical algorithms for offline RL with strong expert-level privacy guarantees, building on top of any gradient-based offline RL algorithm. In contrast with prior approaches which use noisy statistics of the data in offline RL (Qiao and Wang,, 2024), we develop a novel scheme which also identifies a subset of the data which can be used in learning without any added noise. This capability, which we describe in more detail below, is not feasible in the typical trajectory-level privacy settings, and highlights the benefits of systematically studying the expert-level privacy setting.
  • We perform a proof-of-concept evaluation of our algorithm on standard RL benchmarks and compare our approach with an adaptation of DP-SGD to expert-level privacy as a baseline for privatizing the gradient-based offline RL algorithm. We show that our approach significantly improves upon this baseline across multiple test environments. The full results are discussed in Section 6.
@@
Our algorithm: Our algorithm proceeds in two stages. The first is a data filtering stage that produces a set of trajectory prefixes, which can be used in training without any further noise addition, in a privacy-preserving manner. The intuition for this set is that it is quite likely to occur across many experts, and hence using it in training does not violate the privacy of an individual expert. The second part of the algorithm is to run DP-Stochastic Gradient Descent (DP-SGD) [^1] DP-SGD is to an SGD based approach where gradients coming from individual experts are clipped, and appropriate Gaussian noise is added to ensure DP. (Song et al.,, 2013; Bassily et al.,, 2014), but with parameters adapted for expert-level privacy on the remainder of the trajectories. For the first filtering part, we use a variant of sparse vector technique for privacy accounting (Dwork and Roth,, 2014). Once this set of trajectory prefixes is obtained, which we call $D^{\Pi}_{stable}$, for the remainder of the trajectories $D^{\Pi}\setminus D^{\Pi}_{stable}$ we run DP-SGD. Our empirical results strongly show that combining steps one and two of the algorithm described above, performs better than either of them individually. The only assumption we make on the underlying offline RL algorithm is that it is gradient-based, so that we can use DP-SGD to privatize its updates.
@@ 2 RELATED WORK
Comparison to prior work on DP Offline RL: There have been recent work exploring the intersection of differential privacy (DP) and reinforcement learning (RL) (Wang and Hegde,, 2019; Qiao and Wang,, 2023, 2024; Zhou,, 2022; Liao et al.,, 2023)(Qiao and Wang,, 2023, 2024; Zhou,, 2022) explores the problem either in the tabular setting or in the linear function approximation setting. In either of the cases there is a sufficient statistic (like discrete state-action visit counts or feature covariances to privatize sensitive data) which if made DP, would suffice for the whole algorithm to be DP. In our setting, these algorithms are not applicable because we operate in continuous state space, and with general functional approximation for the RL tasks.
Additionally, Wang and Hegde, (2019) explores the use of differential privacy (DP) in continuous spaces to only protect the value function approximator. Liao et al., (2023) explores RL with linear function approximation under local differential privacy guarantee, a stronger notion of privacy where users contributing data do not trust the central aggregator. For reasons mentioned earlier, these works are also not applicable to our setting.
@@
Furthermore, to the best of our knowledge, expert-level DP has not been dealt with in the context of RL so far. Prior work (cited above) are often limited to exploring trajectory-level privacy, which has finer granularity of privacy protection which makes the setting arguably “easier”. In this work, we adopt an expert-level notion of neighbourhood when a large number of behavioural policies (or experts) are used for data collection.
@@
Background on DP-SGD and Sparse-vector: DP-SGD first introduced in (Song et al.,, 2013; Bassily et al.,, 2014), and adapted for deep learning by Abadi et al., (2016), proposed gradient noise injection to guarantee differentially private model training. Several variants of DP-SGD have been proposed such as McMahan et al., (2017), which allows user-level differentially private training in federated learning scenarios and Kairouz et al., (2021) a private variant of the Follow-The-Regularized-Leader algorithm that compares favourably with amplified DP-SGD (privacy amplification by sampling/shuffling (Abadi et al.,, 2016; Bassily et al.,, 2014)).
@@
Sparse-vector algorithm (Roth and Roughgarden,, 2010; Hardt and Rothblum,, 2010; Dwork and Roth,, 2014) develovep over a sequence of works, allows one to carefully track the prviacy budget across queries, and often provide tighter privacy guarantee as opposed to vanilla composition Dwork and Roth, (2014). At a high-level, the idea is to identify “stable queries”, i.e., the queries that do not change their answers by moving to neighboring data set, and not pay the cost of privacy for answering those queries. In Algorithm 2, we use a modification of this idea to output parts of the expert trajectories that are “stable” without incurring the cost of privacy (in $\varepsilon$).
@@
Relationship with the offline RL literature: As mentioned above, our approach is agnostic to the underlying offline RL algorithm used, and can be combined with off-the-shelf model-based(Kidambi et al.,, 2020; Yu et al.,, 2020, 2021) or model-free (Kumar et al.,, 2019, 2020; Cheng et al.,, 2022) offline RL algorithms. We focus on model-free approaches in our experiments.
@@ 3 PRELIMINARIES
We start by introducing the relevant background in offline RL and differential privacy, before formally defining the setting of offline RL with expert-level privacy.
@@ 3.1 Offline RL
Offline RL is a data-driven approach to RL that aims to circumvent the prohibitive cost of interactive data collection in many real-world scenarios. Under this paradigm, a static dataset collected by some behaviour policy $\pi_{\beta}$ is used to train another policy $\pi$ without any further interaction with the environment.
Let the environment be represented by a Markov Decision Process (MDP) $\mathcal{M}=\langle S,A,r,P,\rho_{0},\gamma\rangle$, where $S$ is the state space, $A$ is the action space, $r:S\times A\rightarrow\mathbb{R}$ is the (deterministic) reward function, $P:S\times A\to\Delta(S)$ is the transition kernel ($\Delta(\mathcal{X})$ is the set of all probability distributions over $\mathcal{X}$), $\rho_{0}\in\Delta(S)$ is the initial state distribution and $\gamma\in[0,1]$ is the discount factor.
Definition 1 (Offline RL).
[…]
A key concern in offline RL is that the state-action distribution encountered in the offline data to learn a policy $\pi$ might significantly differ from that encountered upon actually executing $\pi$ in the MDP. There is a host of existing offline RL methods that address this challenge in different ways (see e.g. (Kumar et al.,, 2019; Fujimoto et al.,, 2019; Kumar et al.,, 2020; Cheng et al.,, 2022)). We use CQL (Kumar et al.,, 2020) as our offline RL method due to its strong performance across a variety of tasks in prior evaluations.
We consider a scenario where the dataset contains demonstrations collected from a number of different behavioural policies or experts with varying degrees of optimality. We denote the set of behaviour policies by $\Pi=\{\pi_{1},\dots,\pi_{m}\}$. Note that a requirement of expert-level privacy, which we build towards next, can be in tension with the rewards our learned policy can attain. For instance, if most experts are similar, and one which disagrees from them obtains a very high-reward, then there is an inevitable tradeoff between expert-level privacy and utility.
@@ 3.2 Differential Privacy
Differential privacy (DP) was first introduced in Dwork et al., (2006) to quantify the loss of privacy associated with data released drawn from any statistical database. Informally, a procedure satisfies DP if the distribution of its outputs on two datasets which differ only in one record is very similar. This notion is formally defined below.
Definition 2 ($(\varepsilon,\delta)$-DP).
@@
Definition 2 guarantees that the probability of seeing a specific output on any two neighbouring datasets can differ at most by a factor of $e^{\varepsilon}$, with an additional relaxation of $\delta$. In most ML applications reasonable privacy guarantees are provided by $\varepsilon\leq 10$ (Ponomareva et al.,, 2023) and $\delta=n^{-1.1}$, where $n$ is the number of individuals in the dataset.
One can instantiate Definition 2 with different of neighbourhood relations. Next, we describe the neighbourhood relations underlying our setting of expert-level privacy and the more commonly studied trajectory-level privacy.
Definition 3 (Expert- and Trajectory-level privacy).
Let $\mathcal{D}$ be the domain of all valid trajectories possible in the underlying MDP. Let $D^{\Pi}\in\mathcal{D}^{*}$ be a data set of trajectories generated by a set of experts $\Pi$. We say $D^{\Pi},D^{\Pi^{\prime}}\in\mathcal{D}^{*}$ to be expert-level neighbours if the set of experts $|\Pi\Delta\Pi^{\prime}|=1$, where $\Delta$ is the set difference. Alternatively, if $D,D^{\prime}\in\mathcal{D}^{*}$ are sets of trajectories and $|D\Delta D^{\prime}|=1$, then we say $D$ and $D^{\prime}$ are trajectory-level neighbours.
We note that the key difference between expert and trajectory level privacy notions, as defined above is that removing an expert $\pi$ can remove up to $|D^{\pi}|$ trajectories from the overall dataset, where $D^{\pi}$ is the set of trajectories contributed by $\pi$. This is a much larger change than allowed in trajectory-level privacy, and a naïve treatment would result in paying an additional cost due to group privacy (Dwork and Roth,, 2014), scaling linearly in the number of trajectories each expert contributes. At the same time, as we argue below, via a set of settings, expert-level privacy is often a critically required notion of privacy protection. We also note that we do not include the MDP dynamics and rewards, $P$ and $r$ in the private information we protect, since that would require incorporating changes to the MDP in our neighbourhood relation too.
Example (Personalized healthcare interventions): Many works (Saria,, 2018; Gottesman et al.,, 2019; Tang et al.,, 2022) study the promise of offline RL for learning individual treatment recommendations, using existing clinical data, since real-world data interventions in the form of clinical trials are extremely expensive and limited in their power for learning personalized treatments. Here, the actions comprise of the units of medication administered to an intensive-care patient, now a well-studied setting in the case of sepsis care (Oberst and Sontag,, 2019). States are outcomes of tests as well as relevant information from the patient’s health record including demographic information, recent measaurements of relevant variables such as heart rate, blood pressure etc. The experts correspond to doctors or clinics who choose the observed treatment regimes in the offline samples, and we would like to learn high-performing policies without violating their privacy. Rewards correspond to long-term health outcomes such as recovery time.
As evident in the example above, having expert-level privacy can reduce the barrier for the decision makers to make the decision trajectories available for learning, without worrying about leakage of proprietary information or regulatory/litigation concerns. We note that expert-level privacy differs from the privacy of a patient in the healthcare domains, or broader privacy notions also including the transitions and rewards of the MDP, since we focus on the feasibility of learning from the experts’ choices via offline RL.

### v2

@@ 1 INTRODUCTION
Leveraging existing offline datasets to learn high-quality decision policies via offline Reinforcement Learning (RL) is a critical requirement in many settings ranging from robotics (fu2020d4rl; levine2020offline) and recommendation systems (bottou2013counterfactual; ie2019reinforcement; cai2017real) to healthcare applications (oberst2019counterfactual; tang2022leveraging). Correspondingly, there is now a rich literature on effective techniques (fujimoto2019off; kumar2019stabilizing; kumar2020conservative; cheng2022adversarially) for learning from such offline datasets, collected using one or more behavioural policies. An often overlooked aspect of this literature is, however, that in privacy sensitive domains such as personalized retrieval, advertising, and healthcare, the behavioural policies might reveal private information about the preferences or strategies used by the corresponding experts (users, advertisers, health care providers etc.) whose decisions underlie the offline data. In such scenarios, we ideally seek to uncover the broadly beneficial strategies employed in making decisions by a large cohort of experts whose demonstrations the offline data is collected with, while not leaking the private information of any given expert. We call this novel setting offline RL with expert-level privacy, and design algorithms with strong privacy guarantees for it.
There has been a growing interest in differentially private RL, both in online (wang2019privacy; qiao2023near; zhou2022dprl; liao2023local) and offline (qiao2024offline) settings. The offline RL literature, which is most relevant to this work, primarily focuses on protecting privacy at the trajectory-level. However, when each expert contributes multiple trajectories to the demonstration dataset, private information can still be leaked under trajectory-level differential privacy. Protecting the privacy of an expert in this scenario requires significantly more careful techniques. Secondly, the prior works mostly focus on tabular or linear settings (qiao2023near; qiao2024offline; zhou2022dprl), where the offline RL problem can be effectively solved using count or linear regression based techniques, and is hence amenable to existing techniques for privatizing counts or linear regression models. In contrast, we make no assumptions on the size of the state space, the parameterization of the learned policy or the policy underlying expert demonstrations. For a more detailed discussion of related work, we refer the reader to Section 2.
@@
  • We identify and formalize the problem of offline RL with expert-level privacy and motivate it with practical examples. We also explain the inadequacy of prior algorithms in this setting.
  • We provide practical algorithms for offline RL with strong expert-level privacy guarantees. In contrast with prior approaches which use noisy statistics of the data in offline RL (qiao2024offline) or rely on very strict assumptions like tabular settings or linear function approximators (wang2019privacy), limiting real-world applicability, we remove these constraints, supporting any gradient-based offline RL algorithm with general function approximators and continuous state spaces.
  • DP-SGD (abadi2016dp) is the go-to method today for privately training gradient-based general function approximators. We adapt DP-SGD for expert-level privacy as a natural baseline in our problem setting. Our algorithm identifies a subset of the data which can be used in learning without any added noise, and significantly improves upon the DP-SGD baseline since it selectively adds noise to gradients only where it is needed, unlike DP-SGD which uniformly noises all gradient updates. We evaluate our algorithm on standard RL benchmarks, and discuss the results in Section 6.
Figure 1: Training pipeline for Expert-level Differentially Private Offline RL given an expert set $\Pi=\{\pi_{1},\pi_{2},\dots,\pi_{m}\}$ and input trajectories logged into an offline dataset $D_{\Pi}$. $\mathcal{A}_{DR}$ (Algorithm 2) splits input trajectories and adds stable prefixes to $D^{\Pi}_{stable}$ discarding the rest to $D^{\Pi}_{unst}$. These are used to train any off-the-shelf offline RL algorithm (see Algorithm 3) and learn an expert-level differentially policy $\pi_{private}$.
@@
Our Algorithm: Our algorithm proceeds in two stages; see Figure 1. The first is a data filtering stage that produces a set of trajectory prefixes (denoted $D^{\Pi}_{stable}$), using a variant of the Sparse Vector Technique (SVT) for privacy accounting (dwork2014dp), that can be used in training without any further noise addition, in a privacy-preserving manner. The intuition is that members of this set are likely to occur across many experts, and hence using it in training does not violate the privacy of an individual expert. The second stage involves selectively running DP-Stochastic Gradient Descent (DP-SGD) [^1] DP-SGD is to an SGD based approach where gradients coming from individual experts are clipped, and appropriate Gaussian noise is added to ensure DP. (song2013sgd; bassily2014private; abadi2016dp), adapted for expert-level privacy on the remainder of the trajectories, $D^{\Pi}\setminus D^{\Pi}_{stable}$. Our empirical results strongly show that combining both stages of the algorithm described above, performs better than either of them individually. The only assumption we make on the underlying offline RL algorithm is that it is gradient-based, so that we can use DP-SGD to privatize its updates.
Motivation: We motivate our problem setting using the case of financial portfolio management. Different fund managers employ proprietary and privacy-sensitive investment strategies that guide their trading decisions. In such a setting, treating the algorithms used by individual fund managers as expert policies and learning an expert-level private investment strategy (policy) using offline RL will be beneficial as a starting strategy for new investors and other parties that do not have specialized investment strategies.
Alternatively, in the case of political campaign optimization in a multi-party democracy, large political parties may use proprietary algorithms to maximize their reach, whereas independent candidates, often lacking the same resources, are placed at a disadvantage. Access to a privately learnt aggregate campaigning strategy will allow such players to compete on a somewhat equal footing, allowing voters greater effective choice.
At a high-level, our setting is similar to the ‘warm-start’ problem in optimization (sambharya2024warmstart), translated to an offline RL setting where data is privacy-sensitive and cannot be directly trained on.
Note on Experts: Throughout the paper, we use the term “expert” to refer to RL agents, policies, or algorithms used in various real-world settings. They do not refer to human experts like doctors, trading agents etc. We learn an expert-level DP policy, from the offline trajectories logged by such agents interacting with the environment like behavioural policies in offline RL (levine2020offline; prudencio2023offline).
@@ 2 RELATED WORK
Comparison to Prior Work on Private Offline RL: There have been recent works exploring the intersection of differential privacy (DP) and reinforcement learning (RL) (wang2019privacy; qiao2023near; qiao2024offline; zhou2022dprl; liao2023local)qiao2023near; qiao2024offline; zhou2022dprl; wang2019privacy explore the problem either in the tabular setting or in the linear function approximation setting. In either of the cases there is a sufficient statistic (like discrete state-action visit counts or feature covariances to privatize sensitive data) which if privatized, would suffice to privatize the whole algorithm. Specifically, wang2019privacy explores the use of differential privacy (DP) in continuous spaces to only protect the value function approximator. liao2023local explores RL with linear function approximation under local differential privacy guarantee, a stronger notion of privacy where users contributing data do not trust the central aggregator. Meanwhile, qiao2023near; qiao2024offline; vietri2020private work with tabular MDP settings and discrete state-action spaces, often utilizing some form of state-action visit counts to privatize sensitive data. In our work, we extend existing approaches beyond the state-space and linear function approximator constraints and propose a general algorithm which can be used with continuous state spaces, and with general functional approximation for RL tasks.
@@
To the best of our knowledge, expert-level DP has not been dealt with in the context of RL so far. Prior works often limit themselves to exploring trajectory-level privacy, which has finer granularity of privacy protection which makes the setting arguably “easier". In this work, we adopt an expert-level notion of neighbourhood when a large number of behavioural policies (or experts) are used for data collection. We further note that expert-level privacy differs from existing privacy notions like Joint DP (chowdhury2022dpregret), or other notions which aim to protect the transitions and rewards of the MDP.
@@
Background on DP-SGD and Sparse-Vector Technique: DP-SGD first introduced in song2013sgd; bassily2014private, and adapted for deep learning by abadi2016dp, proposed gradient noise injection to guarantee differentially private model training. Several variants of DP-SGD have been proposed such as mcmahan2017dpfed, which allows user-level DP training in federated learning scenarios and kairouz2021dpftrl a private variant of the Follow-The-Regularized-Leader (DP-FTRL) algorithm that compares favourably with amplified DP-SGD (abadi2016dp; balle2018amplify).
@@
The Sparse-Vector Technique (roth2010interactive; hardt2010multiplicative; dwork2014dp), developed over a sequence of works, allows one to carefully track the privacy budget across queries, and provide tighter data-dependent privacy guarantees as opposed to vanilla composition (dwork2014dp). At a high-level, the idea is to identify “stable queries”, which have low local sensitivity and do not change their answers when moving to a neighboring dataset. Such queries can be answered without paying any additional privacy cost. In Algorithm 2, we use a modification of this idea to output parts of the expert trajectories that are “stable” without incurring any additional privacy cost (in $\varepsilon$).
@@
Relationship with Offline RL Literature: Our approach is agnostic to the underlying offline RL algorithm used and can be combined with any off-the-shelf model-based (kidambi2020morel; yu2020mopo; yu2021combo) or model-free (kumar2019stabilizing; kumar2020conservative; cheng2022adversarially) offline RL algorithms which use gradient updates to learn an optimal policy. We focus on model-free approaches in our experiments.
@@ 3.1 Offline RL
Offline RL (levine2020offline; prudencio2023offline) is a data-driven approach to RL that aims to circumvent the prohibitive cost of interactive data collection in many real-world scenarios. Under this paradigm, a static dataset collected by some (possibly sub-optimal) behaviour policy $\pi_{\beta}$ is used to learn a policy $\pi$ without any interaction with the environment.
Let the environment be modeled by a Markov Decision Process (MDP) $\mathcal{M}=\langle S,A,r,P,\rho,\gamma\rangle$, where $S$ is the state space, $A$ is the action space, $r:S\times A\rightarrow\mathbb{R}$ is the reward function, $P:S\times A\to\Delta(S)$ is the transition kernel ($\Delta(\mathcal{X})$ is the set of all probability distributions over $\mathcal{X}$), $\rho\in\Delta(S)$ is the initial state distribution and $\gamma\in[0,1]$ is the discount factor.
Definition 3.1 (Offline RL).
Given an environment represented by the MDP $\mathcal{M}=\langle S,A,r,P,\rho,\gamma\rangle$ defined above and a dataset $D=\{(s_{i},a_{i},r_{i},s_{i}^{\prime})_{i}\}_{i=1}^{n}$ generated with $(s_{i},a_{i})\sim\mu$ for some state-action distribution $\mu$, $r_{i}=r(s_{i},a_{i})$ and $s_{i}^{\prime}\sim P(\cdot|s_{i},a_{i})$, the objective is to find an optimal policy $\pi^{*}:S\rightarrow\Delta(A)$, defined as
$\pi^{*}=\underset{\pi}{\operatorname*{arg\,max}}\;\mathbb{E}^{\pi}_{s_{1}\sim\rho}[R_{\tau}]$
, where $R_{\tau}=\sum_{t=1}^{|\tau|}\gamma^{t}r_{t}$ is the discounted return of trajectory $\tau$, generated from $\pi$ and $\mathcal{M}$ as: $s_{1}\sim\rho$, $a_{t}\sim\pi(.|s_{t})$, $r_{t}=r(s_{t},a_{t})$ and $s_{t+1}\sim P(.|s_{t},a_{t})$ for $t\geq 1$ upto $|\tau|$ timesteps.
A key concern in offline RL is that the state-action distribution encountered in the offline data to learn a policy $\pi$ might significantly differ from that encountered upon actually executing $\pi$ in the MDP. There is a host of existing offline RL methods that address this challenge in different ways  (kumar2019stabilizing; fujimoto2019off; kumar2020conservative; cheng2022adversarially). We use BCQ (fujimoto2019off) and CQL (kumar2020conservative) as our offline RL methods due to their strong performance across a variety of tasks in prior evaluations and popularity in existing offline RL literature..
We consider a scenario where the dataset contains demonstrations collected from a number of different behavioural policies or experts with varying degrees of optimality. We denote the set of behaviour policies by $\Pi=\{\pi_{1},\dots,\pi_{m}\}$. Imposing the constrain of expert-level privacy can conflict with the rewards attainable by the learned policy. For example, if most experts are similar but a single, high-reward expert diverges from them, an inherent tradeoff arises between expert-level privacy and utility. We only consider the privatization of the expert policies (the chosen actions). Thus, information about the environmnent dynamics such as the transition kernel and reward function are not protected under our proposed algorithm.
@@ 3.2 Differential Privacy
Differential privacy (DP) (dwork2006dp; dwork2014dp) quantifies the loss of privacy associated with data release from any statistical database. Informally, a procedure satisfies DP if the distribution of its outputs on two datasets which differ only in one record is very similar. We use approximate-DP (or $(\varepsilon,\delta)$-DP) throughout our work. This notion is formally defined below.
Definition 3.2 ($(\varepsilon,\delta)$-DP).
@@
Definition 3.2 guarantees that the probability of seeing a specific output on any two neighbouring datasets can differ at most by a factor of $e^{\varepsilon}$, with a relaxation of $\delta$. In most ML applications reasonable privacy guarantees have $\varepsilon\leq 10$ and $\delta\leq\frac{1}{n}$ (ponomareva23dpfy), where $n$ is the cardinality of the dataset.
One can instantiate Definition 3.2 with different of neighbourhood relations. Definition 3.3 describes our setting of expert-level privacy and the more commonly studied trajectory-level privacy.
Definition 3.3 (Expert- and Trajectory-level privacy).
Let $\mathcal{D}$ be the domain of all valid trajectories possible in the underlying MDP. Let $D^{\Pi}\in\mathcal{D}^{*}$ be a data set of trajectories generated by a set of experts $\Pi$. $D^{\Pi},D^{\Pi^{\prime}}\in\mathcal{D}^{*}$ are  expert-level neighbours if the set of experts $|\Pi\Delta\Pi^{\prime}|=1$, where $\Delta$ is the set difference. Alternatively, if $D,D^{\prime}\in\mathcal{D}^{*}$ are sets of trajectories and $|D\Delta D^{\prime}|=1$, then $D$ and $D^{\prime}$ are trajectory-level neighbours.
We note that the key difference between expert and trajectory level privacy notions, as defined above is that removing an expert $\pi$ can remove up to $|D^{\pi}|$ trajectories from the overall dataset, where $D^{\pi}$ is the set of trajectories contributed by $\pi$. This is a much larger change than allowed in trajectory-level privacy, and a naïve treatment would result in paying an additional cost due to group privacy (dwork2014dp), scaling linearly in the number of trajectories each expert contributes. At the same time, as we argue below, via a set of settings, expert-level privacy is often a critically required notion of privacy protection.